Southwest dry cleaners believes that it will need new equipment in 9 years. The equipment will cost $26,000. How much should be invested today at 5% compounded semiannually, to yield $26,000?
Answer by stanbon(75887)

Southwest dry cleaners believes that it will need new equipment in 9 years. The equipment will cost $26,000. How much should be invested today at 5% compounded semiannually, to yield $26,000?
----------------
A(t) = P(1+(r/n))^(nt)
-----
26000 = P(1+(0.05/2))^(2*9)
-------
26000 = P(1 + 0.025)^18
P = 26000/(1.5597)
P = $16670.31
------------------
Cheers,
stan H.